OpenStudy (anonymous):

solve for x, show work -14 = -4 (9x - 1)

OpenStudy (anonymous):

where you dont undestand this question??

OpenStudy (anonymous):

after distributing -4 -14 = -36x + (-4) -4 - (-14) = -36x + (-4) - (-4) -18 = - 36x and then i would simplify, but it doesnt seem right. i dont understand where i made a mistake

OpenStudy (anonymous):

ok ok here is the mistake

OpenStudy (anonymous):

after distributing -4 -14 = -36x+4 now do it

OpenStudy (anonymous):

if you get x= 1/2 then you are right else wrong

OpenStudy (anonymous):

4 - (-14) = -36x + 4 - 4 18 = 36x simplified its .5?

OpenStudy (anonymous):

yes .5 = 1/2 you got it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome